Joe has a collection of nickels and dimes that is worth $2.10. If the number of dimes was tripled and the number of nickels was increased by 24, the value of the coins would be $5.70. How many nickels and dimes does he have?
Answer by lwsshak3(11628)

Joe has a collection of nickels and dimes that is worth $2.10. If the number of dimes was tripled and the number of nickels was increased by 24, the value of the coins would be $5.70. How many nickels and dimes does he have?
**
let x=Number of nickels
let y=Number of dimes
..
.05x+.10y=2.10
.05(x+24)+3*.10=5.70
..
.05x+.10y=2.10
.05x+1.2+.30y=5.70
..
.05x+.10y=2.10
.05x+.30y=4.50
subtract
-.2y=-2.4
y=12
..
.05x+.10y=2.10
.05x+1.2=2.1
.05x=2.1-1.2=.90
x=.90/.05
x=18
Number of nickels=18
Number of dimes=12